1. Let f be the function deÖned by

f (x) = x^2/(-2x + 1)^2

(a) Determine the vertical and horizontal asymptotes (show all limits).

(b) Use the sign pattern for f'(x) to determine

(i) the interval(s) over which f rises and where it falls;

(ii) the local extrema.

(c) Use the sign pattern for f''(x) to determine

(i) where the graph of f is concave up and where it is concave down.

(ii) the inflection points (if any)


1
Expert's answer
2022-02-21T16:17:59-0500

(a)


"f (x) =\dfrac{x^2 }{(-2x+1)^2}"

"-2x+1\not=0=>x\not=1/2"

"D(f): (-\infin, 1/2)\cup (1/2 , \infin)"


"\lim\limits_{x\to(1/2)^-}\dfrac{x^2 }{(-2x+1)^2} =\infin"

"\lim\limits_{x\to(1/2)^+}\dfrac{x^2 }{(-2x+1)^2} =\infin"

Vertical asymptote: "x=1\/2"


"\lim\limits_{x\to-\infin}\dfrac{x^2 }{(-2x+1)^2} =\lim\limits_{x\to-\infin}\dfrac{x^2/x^2 }{(-2x/x+1/x)^2}"

"=\lim\limits_{x\to-\infin}\dfrac{1}{(-2+1/x)^2}=1/4"


"\lim\limits_{x\to\infin}\dfrac{x^2 }{(-2x+1)^2} =\lim\limits_{x\to\infin}\dfrac{x^2/x^2 }{(-2x/x+1/x)^2}"

"=\lim\limits_{x\to\infin}\dfrac{1}{(-2+1/x)^2}=1/4"

Horizontal asymptote: "y=1\/4"

There is no slant (oblique) asymptote.


(b)


"f'(x)=(\dfrac{x^2 }{(-2x+1)^2} )'"

"=\dfrac{2x(-2x+1)^2-x^2(2)(-2)(-2x+1) }{(-2x+1)^4}"

"=\dfrac{-4x^2+2x+4x^2}{(-2x+1)^3} =\dfrac{2x}{(-2x+1)^3}"

Find critical number(s)


"f'(x)=0=>\dfrac{2x}{(-2x+1)^3} =0"

"x=0"

(i)

If "x<0, f'(x)<0, f(x)" decreases.

If "0<x<1\/2, f'(x)>0, f(x)" increases.

If "x>1\/2, f'(x)<0, f(x)" decreases.

The function "f" increases on "(0, 1\/2)."

The function "f" decreases on "(-\infin, 0)\cup (1/2, \infin)."

(ii)

The function "f" has a local minimum at "x=0."

The function "f" has no a local maximum.


(c)


"f''(x)=(\dfrac{2x }{(-2x+1)^3} )'"

"=2\cdot\dfrac{(-2x+1)^3-x(3)(-2)(-2x+1)^2 }{(-2x+1)^6}"

"=2\cdot\dfrac{-2x+1+6x}{(-2x+1)^4} =\dfrac{8x+2}{(-2x+1)^4}"

Find the inflection point(s)


"f''(x)=0=>\dfrac{8x+2}{(-2x+1)^4} =0"

"x=-1\/4"

(i)

If "x<-1\/4, f''(x)<0, f(x)" is concave down.

If "-1\/4<x<1\/2, f''(x)>0, f(x)" is concave up.

If "x>1\/2, f''(x)>0, f(x)" is concave up.

The function "f" is concave up on "(-1/4, 1/2)\cup (1/2, \infin)."

The function "f" is concave down on "(-\infin, -1/4)."

(ii)

The function "f" has inflection point at "x=-1\/4."